OpenStudy (anonymous):

Verify that each equation is an identity 1+sinA-cos2A ------------- = tanA cosA+sin2A

OpenStudy (anonymous):

anyone?

OpenStudy (cwrw238):

1 + sinA - cos2A = 1 + sinA - (1 - 2sin^2A) = 2sin^2A + sin A = sinA(2sinA + 1) cosA + sin 2A = cosA + 2sinAcosA = cosA(2sinA + 1) sinA(2sinA + 1) ------------- = sinA / cosA = tan A cosA(2sinA + 1) this proves the identity

OpenStudy (anonymous):

thank you!!!!

OpenStudy (cwrw238):

yw